Educare, a nonprofit, teaches elementary students after school. In five years it plans to replace the current vehicle used to chauffer children from their respective schools to the Educare facility. Educare expects to pay $30,000 for the new van. If it can invest money at 5% compounded quarterly, how much must it invest today?"
Answer #1

Rate = 5% / 4 = 1.25%

Number of periods = 5 * 4 = 20

Present value = FV / (1 + r)n

Present value = 30,000 / (1 + 0.0125)20

Present value = 30,000 / 1.282037

Present value = $23,400.26

It must invest $23,400.26